unseen:

Given: y varies directly with x and has a constant rate of change of 7. When the value of y is 12, then the value of x would be _____. equal to zero less than 2 equal to 7 greater than 12

unseen:

@Falconmaster

unseen:

tagged yah

unseen:

hey @eatsalad how you

Falconmaster:

im thinking c

unseen:

im thinking im bad at math lol

unseen:

eat wanna answer if you have a different answer and ill see if yall are right or wrong

unseen:

brb dont leave

unseen:

all right im back but im getting the answer from my teacher to see if we are right she said hold on a minute

unseen:

so i will post the next question in this question or well in the reply bar

unseen:

What is the equation of the line, in slope-intercept form, that passes through (3, -1) and (-1, 5)? y = - 3/2x +7/2 <---- fraction 2y = -3x + 7 2x + 3y - 7 = 0

eatsalad:

question one Y/x=7 12x=7 7x=12 x=12/7=1 5/7 So now what do you think the answer would be?

unseen:

Hmmm c

eatsalad:

No

unseen:

b

eatsalad:

Yes, nice job.

unseen:

yeah im 1% smarter

unseen:

lol thanks eat